# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ Appendix F Liver Fairness
The US liver transplant system has seen several key policy changes designed to optimize organ distribution and improve patient outcomes. In 2002, the model for end-stage liver disease (MELD) was adopted to determine patients’ priority for liver transplants based on the severity of their illness [51]. In June 2013, UNOS implemented the Share 35 policy to require that the regional sharing of livers for patients with a MELD score of $\geq 35$ prioritizes the local sharing of livers to less severe patients with a MELD score of $<35$ [94]. In January 2016, the MELD-sodium (i.e. MELD-Na) score went into effect to improve the original MELD by incorporating serum sodium levels into the calculation [88]. In February 2020, the acuity-circle (AC) policy for distributing livers replaced the longstanding donation service area based policy [86]. The AC policy uses a series of concentric circles, centered on the donor hospital, to determine the distribution range for available livers. In July 2023, the calculation formula for MELD was updated to include additional variables including female sex and serum albumin [87]. The new model, known as MELD 3.0, was expected to further refine the sorting of transplant candidates based on their medical urgency for transplant [41].
Feature Selection
We use the following features in the case study. Except for the liver allocation outcome, all features are determined at time of registration or listing. Thus (A1) holds. These features include all the possible parents to the liver allocation outcome available in the OPTN dataset, thus satisfying (A2). For time period 2017-2019, total observations is $n=21,101$ patients. For time period 2020-2022, total observations is $n=22,807$ patients. Below, we provide the definition of each variable included in our causal fairness analysis. Summary statistics are reported in Table F.1.
@@
Table F.1: Mean values (standard deviations) for features with statistically significant differences between males and females ($\alpha=0.05$). Summary statistics for all features are available on GitHub (https://anonymous.4open.science/r/LD3-4440).

### v3

@@ E.2 Liver Transplant Allocation
The US liver transplant system has seen several key policy changes designed to optimize organ distribution and improve patient outcomes. In 2002, the model for end-stage liver disease (MELD) was adopted to determine patients’ priority for liver transplants based on the severity of their illness (Malinchoc et al. 2000). In June 2013, UNOS implemented the Share 35 policy to increase access to organs for patients with MELD scores $\geq 35$, given the high waitlist mortality in this subgroup (Wald et al. 2013). In January 2016, the MELD-sodium (i.e., MELD-Na) score incorporated serum sodium levels into the calculation (UNOS 2015). In February 2020, the acuity-circle (AC) policy for distributing livers replaced the longstanding donation service area-based policy (UNOS 2020). The AC policy uses a series of concentric circles, centered on the donor hospital, to determine the distribution range for available livers. In July 2023, the calculation formula for MELD was updated to include additional variables, including female sex and serum albumin (UNOS 2022). The new model, known as MELD 3.0, was expected to further refine the sorting of transplant candidates based on their medical urgency for transplant (Kim et al. 2021).
Data Preprocessing and Feature Selection
We used the National Standard Transplant Analysis and Research (STAR) dataset for patient-level information on waitlisted transplant candidates, recipients, and donors from the Organ Procurement & Transplantation Network (OPTN) up to October 2022 (OPTN 2024). We filtered to include adult patients awaiting or having undergone a whole liver transplant from deceased donors in 2017-2019 and 2020-2022, keeping features relevant for discrimination discovery (Appendix E.2). Time frames were selected as stable periods without changes in the UNOS allocation policy.
We used the following features in the case study. Except for the liver allocation outcome, all features were determined at time of registration or listing. Thus, A1 holds. These features include all the possible parents to the liver allocation outcome available in the OPTN dataset, thus satisfying A2 to the best of our knowledge. For time period 2017-2019, total observations was $n=21,101$ patients. For time period 2020-2022, total observations was $n=22,807$ patients. Below, we provide the definition of each variable included in our causal analysis. Summary statistics are reported in Table E.6.
@@
Unos Policy (2017–2019) Unos Policy (2020–2022) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ active exception case, diagnosis, education, ethnicity, exception type, initial age, initial MELD, region, weight active exception case, diagnosis, education, ethnicity, exception type, initial age, initial MELD, recipient blood type, region, weight active exception case, diagnosis, education, ethnicity, initial age, initial MELD, payment method, region, weight active exception case, diagnosis, education, ethnicity, initial age, initial MELD, payment method, region, weight SDC 1 1 1 1 WCDE 0.033 [0.016,0.049] 0.025 [0.008,0.042] 0.032 [0.018,0.047] 0.032 [0.018,0.047] $p$-value 0.000 0.003 0.000 0.000
Table E.2: Results obtained with nonparametric $\chi^{2}$ independence tests and double machine learning. Total features is 18 (Appendix E.2). For time period 2017-2019, total observations is $n=21101$ patients. For time period 2020-2022, total observations is $n=22807$ patients. With no covariate control, SDC = 1 and the weighted CDE is 0.049 [0.033,0.066] ($p$-value = 0.000).
Table E.3: SDC assessment and weighted CDE estimation with double machine learning using results from PC Algorithm (Spirtes, Glymour, and Scheines 2001), as implemented by Kalisch and Bühlman (2007). We define $\mathbf{A}_{\text{{DE}}}$ using the union method (PC; Appendix D.2), though results for $\mathbf{A}_{\text{{DE}}}$ were exactly the same using PC.
Unos Policy (2017–2019) Unos Policy (2020–2022) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ payment method payment method, region active exception case, diagnosis, initial age, initial MELD, recipient blood type diagnosis, recipient blood type, region SDC 1 1 1 1 WCDE 0.049 [0.032,0.065] 0.048 [0.032,0.064] 0.032 [0.019,0.046] 0.032 [0.017,0.048] $p$-value 0.000 0.000 0.000 0.000
Table E.4: SDC assessment and weighted CDE estimation with double machine learning using results from LDECC using the union method for selection $\mathbf{A}_{\text{{DE}}}$ (LDECC; Appendix D.2) (Gupta, Childers, and Lipton 2023). Results for $\mathbf{A}_{\text{{DE}}}$ were almost concordant using LDECC, with the exception of 2020-2022 when $\alpha=0.05$ (where the intersection method excluded diagnosis).
Table E.5: Total number of independence tests performed and runtimes in seconds for LD3 vs PC (Kalisch and Bühlman 2007) and LDECC (Gupta, Childers, and Lipton 2023) on the liver transplant allocation dataset. While PC is worst-case exponential in time complexity (Spirtes, Glymour, and Scheines 2001), LDECC is worst-case exponential for some structures and polynomial for others (Gupta, Childers, and Lipton 2023). This could explain the wide variation in LDECC runtimes across experimental settings.
@@
Table E.6: Mean values (standard deviations) for features with statistically significant differences between males and females ($\alpha=0.05$). Summary statistics for all features are available on GitHub (https://anonymous.4open.science/r/LD3-4440).